OpenStudy (anonymous):

show all work. Find the difference: (6x2 + x + 7) – (6x2 – 4x – 5)

OpenStudy (inkyvoyd):

We aren't here to do your homework for you.

OpenStudy (inkyvoyd):

Please only ask a question if you are seeking assistence in learning how to solve it.

OpenStudy (anonymous):

am not asking you to do my home work thank you i dont understand this at all math is a very hard subject for me

OpenStudy (inkyvoyd):

Alright, that's ok.

OpenStudy (inkyvoyd):

multiply the stuff in the minus sign by negative one.

OpenStudy (anonymous):

Combine like terms, make sure to keep track of the negative signs (a negative applied to a negative is the same as a positive).

OpenStudy (inkyvoyd):

"– (6x2 – 4x – 5)"=6x^2+4x+5 Combine like terms.

OpenStudy (anonymous):

would i combine like term next

OpenStudy (anonymous):

The negative distributes to the 6x^2 term as well (it's negative).

OpenStudy (inkyvoyd):

whenever you see a minus in front of parenthesis, it means to subtract what in the parenthesis, or, add the negative of everything in the parenthesis. for example (x^2-3x+4)-(2x^2+2x-5) =x^2-3x+4-2x^2-2x+5 =-x^2-5x+9 In fact don't be scared by the variables (the x). Treat it just like you would any number. Let's pretend that x is 5. (5^2-3*5+4)-(2*5^2+2*5-5) how would you simplify it?

OpenStudy (anonymous):

sorry am still lost

OpenStudy (anonymous):

that didnt really help

OpenStudy (anonymous):

some one help

OpenStudy (saifoo.khan):

Where are u stuck?

OpenStudy (anonymous):

 combine like terms - what do you get?